### v1

# 93D15, 93C10, 93B51, 93D30

Abstract
We consider the stabilization problem for driftless control-affine systems under the bracket-generating condition. In our previous works, a class of time-varying feedback laws has been constructed to stabilize the equilibrium of a nonholonomic system under rather general controllability assumptions. This stabilization scheme is based on the sampling concept, which is not equivalent to the classical definition of solutions for the corresponding nonautonomous closed-loop system. In the present paper, we refine the previous results by presenting sufficient conditions for the convergence of classical solutions of the closed-loop system to the equilibrium. Our theoretical findings are applied to a multidimensional driftless control-affine system and illustrated through numerical simulations.
keywords:
nonlinear control, nonholonomic system, stabilization, time-varying feedback law, Chen–Fliess series
1 INTRODUCTION
Consider a driftless control-affine system
$\dot{x}=\sum_{j=1}^{m}u_{j}f_{j}(x),\quad x\in D\subset{\mathbb{R}}^{n},\;u\in{\mathbb{R}}^{m},\;m<n,$ (1.1)
with the state $x=(x_{1},x_{2},...,x_{n})^{\top}$ and the control $u=(u_{1},u_{2},...,u_{m})^{\top}$. We assume that the vector fields $f_{1}$, …, $f_{m}$ are smooth in the domain $D$, with $0\in D$, and that
${\mathrm{rank}}\left(f_{1}(0),f_{2}(0),...,f_{m}(0)\right)=m.$
This class of systems includes, in particular, mathematical models of underactuated robots with nonholonomic constraints. It is well-known that the equilibrium $x=0$ of the underactuated system (1.1) is not stabilizable by a regular time-invariant feedback law $u=h(x)$ due to the violation of Brockett’s condition [1].
The stabilization problem for underactuated systems under nonlinear controllability conditions has been extensively studied. Without pretending to be exhaustive, we mention stabilizability results for general controllable systems using time-invariant discontinuous controllers [2] and time-varying feedback laws [3][4, Chap. 11], as well as averaging techniques with fast oscillating inputs [5] and control design with transverse functions [6, 7].
In [8], a class of nonholonomic systems with external disturbances is considered, and constrained stabilizing feedback laws are proposed using state transformations and dynamic surface control. These controllers are experimentally tested on a wheeled mobile robot. A chained nonholonomic system in cascaded form with two subsystems is considered in [9], accounting for model uncertainties and disturbances. The stabilization problem is addressed using a state feedback law, where output control is applied to the fully actuated subsystem, and an exponentially stable component is designed. In [10], a fuzzy logic approach is applied to approximate uncertainties in a kinematic nonholonomic system to address the stabilization problem with output constraints. The considered constraints are satisfied by combining the design of a barrier Lyapunov function with an event-triggered scheme, which is applied to a wheeled mobile robot.
In the paper [11], a family of time-varying feedback laws was proposed to solve the exponential stabilization problem fo nonholonomic systems under the step-2 bracket-generating condition. The coefficients of these controllers are defined by solving an auxiliary system of nonlinear algebraic equations, and a simplified control design procedure was developed in [12] using the inverse of the matrix that appears in the controllability rank condition. Additionally, an algorithm for constructing stabilizing controllers for high-order nonholonomic systems was recently reported in [13]. In the above-mentioned papers [12, 11, 13], the solutions of the closed-loop system are defined in the sense of sampling, leaving open the question of the asymptotic behavior of classical (or Carathéodory) solutions under the proposed Lie-bracket-based oscillating feedback laws. The present work aims to fill that gap by demonstrating that time-varying controllers, whose oscillating components are designed similarly to those in [12, 11], ensure the convergence of classical solutions to equilibrium.
The subsequent exposition is organized as follows. Section 2 presents our control design scheme, which is based on a weak control Lyapunov function and establish strict decrease conditions using suitable oscillating controllers. The resulting convergence results are formulated in Theorem 2.1, while Theorem 2.2 extends the design by incorporating our previous sampling-based strategies. In Section 3, this control design methodology is applied to a multidimensional nonholonomic integrator. Finally, concluding remarks are provided in Section 4.
2 CONTROL DESIGN
Let us consider a positive definite function $V\in C^{2}(D)$ and denote by
$\dot{V}(x,u)=\sum\limits_{j=1}^{m}u_{j}L_{f_{j}}V(x)$
the time derivative of $V(x)$ along the trajectories of (1.1), where $L_{f_{j}}V(x):=\sum\limits_{i=1}^{n}\frac{\partial V(x)}{\partial x_{i}}f_{ji}(x)=\nabla V(x)f_{j}(x)$ stands for the directional derivative of $V(x)$ along $f_{j}(x)$ (the gradient $\nabla V(x)=\Big{(}\dfrac{\partial V}{\partial x_{1}},...,\dfrac{\partial V}{\partial x_{n}}\Big{)}$ is treated as row vector in these notations). Note that any such $V(x)$ is a weak control Lyapunov function (weak CLF) because of the driftless form of system (1.1), i.e. for each $x\in D$, there exists a $u_{x}\in{\mathbb{R}}^{m}$ such that $\dot{V}(x,u_{x})\leq 0$. Note that there are no strict control Lyapunov functions $V(x)$ for system (1.1) under our assumptions because of Brockett’s [1] and Artstein’s [14] theorems (see also [15]). It means that $\dot{V}(x,u)$ cannot be made negative definite by any choice of controls. Thus, for any positive definite $V(x)$ and any neighborhood of the origin in $\mathbb{R}^{n}$, there are points $x\neq 0$ with $\dot{V}(x,u)\geq 0$ for all $u\in{\mathbb{R}}^{m}$.
We propose a solution of the asymptotic stabilization problem by constructing a family of time-varying feedback controls $u=h^{\varepsilon}(x,t)$ depending on a small parameter $\varepsilon>0$. Even though the time derivative of $V(x)$ along the classical solutions $x(t)$ of the corresponding closed-loop system is not negative definite, we will derive sufficient conditions for the finite difference $V(x(t_{0}+\varepsilon))-V(x(t_{0}))$ to be negative along nontrivial solutions, provided that $\varepsilon$ is small enough. Then, under additional technical assumptions, the above property implies the convergence of $x(t)$ to 00 as $t\to+\infty$.
For the sake of brevity of presentation, we restrict ourselves to a particular case of step-2 bracket-generating systems. Namely, let $S\subset\{1,2,...,m\}^{2}$ be a set of pairs of indices $I=(ij)$ such that $|S|=n-m$ and
${\mathrm{span}}\{f_{k}(x),f^{I}(x)|k=\overline{1,m},\,I\in S\}={\mathbb{R}}^{n}\;\text{for all}\;x\in D,$
where $f^{I}(x):=[f_{i},f_{j}](x)$ is the Lie bracket of vector fields $f_{i}(x)$ and $f_{j}(x)$ with the indices $(ij)=I$:
$[f_{i},f_{j}](x):=\dfrac{\partial f_{j}(x)}{\partial x}f_{i}(x)-\dfrac{\partial f_{i}(x)}{\partial x}f_{j}(x)$.
Then we construct the time-varying feedback control $u=h^{\varepsilon}(x,t)$ as
$h_{k}^{\varepsilon}(x,t)=v_{k}^{0}(x)+\tilde{\gamma}\sum_{I\in S}v_{k}^{I}(x)\phi_{k}^{I,\varepsilon}(t),\quad k=1,2,...,m,$ (2.1)
where the time-independent part $v^{0}(x)$ is designed from the derivative $\dot{V}(x,u)$ of a weak CLF candidate $V(x)$, $\gamma>0$ is a gain parameter, and the time-varying part $v^{I}(x)\phi^{I,\varepsilon}(t)$ approximately drives the system along the Lie bracket ±$\pm f^{I}(x)$ to ensure the decreasing of $V(x(t))$ with respect to discrete $\varepsilon$-time shifts even if $\dot{V}$ is not negative definite. The functions $\phi_{k}^{I,\varepsilon}(t)$ are defined in the following way: if $I=(ij)\in S$, then
$\phi_{i}^{I,\varepsilon}(t){=}2\sqrt{\frac{\varkappa_{I}\pi}{\varepsilon}}\cos(\varkappa_{I}\omega t),\phi_{j}^{I,\varepsilon}(t){=}2\sqrt{\frac{\varkappa_{I}\pi}{\varepsilon}}\sin(\varkappa_{I}\omega t),$ (2.2)
$\phi_{k}^{I,\varepsilon}(t)=0\;\text{for all}\;k\notin\{i,j\},$
where $\omega=\frac{2\pi}{\varepsilon}$ and $\varkappa_{I}$ is a positive integer frequency multiplier. The latter integer multipliers should be chosen to avoid resonances between the control components that generate different Lie brackets, i.e.
$\varkappa_{I}\neq\varkappa_{I^{\prime}}\quad\text{for all}\;\;S\ni I\neq I^{\prime}\in S.$ (2.3)
In particular, we can take $\varkappa_{I_{1}}=1$, $\varkappa_{I_{2}}=2$, …, $\varkappa_{I_{n-m}}=n-m$ for $S=\{I_{1},I_{2},...,I_{n-m}\}$. The closed-loop system (1.1) with controls $u_{k}=h^{\varepsilon}_{k}(x,t)$ of the form (2.1) reads as
$\dot{x}=g_{0}(x)+\gamma\sum_{I\in S}\sum_{k=1}^{m}g_{k}^{I}(x)\phi_{k}^{I,\varepsilon}(t)$ (2.4)
$\text{with }\quad g_{0}(x)=\sum_{k=1}^{m}v_{k}^{0}(x)f_{k}(x),\;g_{k}^{I}(x)=v_{k}^{I}(x)f_{k}(x),$ (2.5)
where, similarly to (2.2), we assume that, for each $I=(ij)\in S$: $v_{k}^{I}(x)=0\;\text{and}\;g_{k}^{I}(x)=0\;\text{for all}\;k\notin\{i,j\}$. If the vector fields $g_{0}$, $g_{k}^{I}$ are of class $C^{2}$ and the functions $\phi_{k}^{I,\varepsilon}(t)$ are defined by (2.2), then the Chen–Fliess expansion of the solution $x(t)\in D$ of (2.4) with the initial data $x(0)=x^{0}\in D$ can be expressed for small $\varepsilon>0$ as follows [16, 12]:
$x(\varepsilon)=x^{0}+\varepsilon g_{0}(x^{0})+\frac{\varepsilon^{2}}{2}L_{g_{0}}g_{0}(x^{0})$ (2.6)
$\qquad-\varepsilon^{3/2}{\gamma}\sqrt{\pi}\sum_{I\in S}\frac{1}{\sqrt{\varkappa_{I}}}[g_{0}^{I},g_{j}^{I}](x^{0})$
$\qquad+\varepsilon{\gamma}^{2}\sum_{I\in S}[g_{i}^{I},g_{j}^{I}](x^{0})+O(\varepsilon^{3/2})$
$=x^{0}+\varepsilon\bigl{(}g_{0}(x^{0})+{\gamma}^{2}\sum_{I\in S}[g_{i}^{I},g_{j}^{I}](x^{0})\bigr{)}+\varepsilon^{3/2}R(x^{0},\varepsilon,\gamma).$
For each compact $D^{\prime}\subset D$ and each $\gamma>0$, there is a $\tilde{\varepsilon}>0$ such that $\|R(x^{0},\varepsilon,\gamma)\|\leq C_{R}$ for all $(x^{0},\varepsilon)\in\tilde{D}\times(0,\tilde{\varepsilon}]$, where the constant $C_{R}$ can be obtained from Lemma 5.1 of [16]. The Lie brackets appearing in (2.6) can also be represented in terms of the original vector fields $f_{i}$, $f_{j}$ with $(ij)=I$:
$[g_{i}^{I},g_{j}^{I}]=v_{i}^{I}v_{j}^{I}[f_{i},f_{j}]+v_{i}^{I}f_{j}\nabla v_{j}^{I}f_{i}-v_{j}^{I}f_{i}\nabla v_{i}^{I}f_{j},$ (2.7)
under the convention that the gradient $\nabla v_{j}^{I}(x)$ is a row vector. Formula (2.6) allows to express the increment of $V(x(t))$ along the trajectories of (2.4):
$\frac{V(x(\varepsilon))-V(x^{0})}{\varepsilon}=W(x^{0})+\varepsilon^{1/2}r(x^{0},\varepsilon,\gamma),$ (2.8)
where
$W(x)=L_{g_{0}}V(x)+\gamma^{2}\sum_{I=(ij)\in S}L_{[g_{i}^{I},g_{j}^{I}]}V(x),$ (2.9)
and $r(x^{0},\varepsilon,\gamma)$ is a bounded remainder which can be estimated in terms of $R(x,\varepsilon,\gamma)$ and the derivatives of $V(x)$. A key idea of our control design is to define the functions $v^{I}_{k}(x)$ and gain $\gamma$ to ensure that $W(x)$ is negative definite. The main result in this area is summarized below.
Theorem 2.1.
Let the function $V(x)$ and vector fields $g_{0}(x)$, $g^{I}_{k}(x)$, $k=\overline{1,m}$, $I\in S$, be twice continuously differentiable in $D$, $\gamma>$ be a positive constant, $D^{\prime}\subseteq D$ be a closed bounded domain such that $0\in{\mathrm{int}}\,D^{\prime}$, and let the following conditions be satisfied:
  • i)
    $V(x)$ is positive definite in $D^{\prime}$;
  • ii)
    $W(x)$ given by (2.9) is negative definite in $D^{\prime}$;
  • iii)
    there exist an $\bar{\varepsilon}>0$ and a $c_{r}>0$ such that
    $|r(x,\varepsilon,\gamma)|\leq-c_{r}W(x)\;\;\text{for all}\;x\in D^{\prime},\;\varepsilon\in(0,\bar{\varepsilon}];$
  • iv)
    $g_{0}(0)=0$ and $g^{I}_{k}(0)=0$ for all $k=\overline{1,m}$, $I\in S$.
Then there exists a $\delta>0$ such that each solution $x(t)$ of system (2.4) with the initial condition $x(0)\in B_{\delta}(0)\subset D^{\prime}$ tends to zero as $t\to+\infty$, provided that the time-varying functions $\phi_{k}^{I,\varepsilon}$ are given by (2.2) under assumption (2.3), and the parameter $\varepsilon>0$ is small enough.
Proof.
Given a function $V(x)$ satisfying condition i), denote its level set as $\mathcal{L}_{c}=\{x\in D:V(x)\leq c\}$ for $c\geq 0$. Let $d>0$ and $\delta>0$ be such that
$\overline{B_{\delta}(0)}\subset\mathcal{L}_{d}\subset{\mathrm{int}}\,D^{\prime}.$
Let $\varepsilon_{0}\in(0,\bar{\varepsilon}]$ be small enough to guarantee that the solutions of system (2.4) with $x(0)=x^{0}\in\overline{B_{\delta}(0)}$ are well-defined in $D^{\prime}$ on the time interval $t\in[0,\varepsilon_{0}]$. Throughout the proof, we assume $\varepsilon\in(0,\varepsilon_{0}]$.
The first step of the proof involves ensuring that the solutions of system (2.4) with $x^{0}\in\mathcal{L}_{d}$ and a sufficiently small $\varepsilon$ evolve in $D^{\prime}$ for $t\in[0,\varepsilon]$, and obtaining some a priori estimates. Denote
$L_{0}{=}\sup_{x\in D^{\prime}}\left\|\frac{\partial g_{0}(x)}{\partial x}\right\|,\,L_{g}{=}\max_{I\in S,k\in\{1,\dots,m\}}\sup_{x\in D^{\prime}}\left\|\frac{\partial g_{k}^{I}(x)}{\partial x}\right\|,$
then, taking into account $g_{0}(0)=g_{k}^{I}(0)=0$ in condition iv), we obtain that
$\|g_{0}(x)\|\leq L_{0}\|x\|,\,\|g_{k}^{I}(x)\|\leq L_{g}\|x\|,$ (2.10)
for all $x\in D^{\prime}$, $I\in S$, $k\in\{1,\dots,m\}$. Furthermore, note that
$\max_{t\in[0,\varepsilon]}\sum_{I\in S}\sum_{k=1}^{m}|\phi_{k}^{I,\varepsilon}(t)|\leq\frac{c_{\phi}}{\sqrt{\varepsilon}}$
with $c_{\phi}=2\sqrt{2\pi}\sum_{I\in S}\sqrt{\varkappa_{I}}$. From the integral representation of solutions of system (2.4),
[…]
Thus, for any $t\in[0,\varepsilon]$,
[…]
Then the Grönwall–Bellman inequality implies
$\|x(t)\|\leq c_{x}\|x^{0}\|\text{ for all }t\in[0,\varepsilon],$ (2.11)
where $c_{x}=e^{L_{0}\varepsilon+L_{g}c_{\phi}\sqrt{\varepsilon}}$. Taking $\varepsilon_{1}>0$ as the minimum of $\varepsilon_{0}$ and the positive solution of the equation $L_{0}\varepsilon+L_{g}c_{\phi}\sqrt{\varepsilon}=\ln\frac{\sup_{x\in D^{\prime}}\|x\|}{\sup_{x\in\mathcal{L}_{d}}\|x\|}$, we conclude that the solutions of system (2.4) with $x^{0}\in\mathcal{L}_{d}$ and $\varepsilon\in(0,\varepsilon_{1}]$ stay in $D^{\prime}$ for all $t\in[0,\varepsilon]$.
The next step is to show that there exists an $\varepsilon_{2}\in(0,\varepsilon_{1}]$ such that the function $V$ satisfies the property $V(x(\varepsilon))<V(x^{0})$ along the solutions of system (2.4) with $\varepsilon\in(0,\varepsilon_{2}]$ For this purpose, we exploit representation (2.8) and assumptions ii)–iii). Namely, let the constants $\bar{\varepsilon}$ and $c_{r}>0$ be defined from the assumptions ii) and iii). Then representation (2.8) implies
$V(x(\varepsilon))\leq V(x^{0})+\varepsilon W(x^{0})(1-\sqrt{\varepsilon}c_{r}).$
Putting $\varepsilon_{2}={\mathrm{min}}\{c_{r}^{2},\varepsilon_{1}\}$, we ensure that $V(x(\varepsilon))<V(x^{0})$ for any $x^{0}\in\mathcal{L}_{d}\setminus\{0\}$ and any $\varepsilon\in(0,\varepsilon_{2}]$. In particular, $x(\varepsilon)\in\mathcal{L}_{d}$.
Repeating the above steps for $x(\varepsilon)$, $x(2\varepsilon)$, …, we conclude that $x(t)\in D^{\prime}$ for all $t\in[0,\infty)$ and obtain the monotonically decreasing sequence $\{V(x(j\varepsilon))\}_{j\in\mathbb{N}}$ with $\inf_{j}V(x(j\varepsilon))=0$. This implies $\lim_{j\to\infty}V(x(j\varepsilon))=0$, and therefore, $\lim_{j\to\infty}\|x(j\varepsilon)\|=0$ because of the continuity and positive definiteness of $V$. Together with (2.11), this yields $\lim_{t\to\infty}\|x(t)\|=0$.
The next theorem justifies the applicability of the controls, proposed previously in papers [11, 16, 12, 13] in the context of sampling, for stabilizing the classical solutions.
Theorem 2.2.
Given a positive definite function $V\in C^{2}(D)$ such that $\nabla V(x)=0$ only at $x=0$, let the functions $v^{0}_{k},v^{I}_{k}\in C^{2}(D)$ be defined by the following formulas for $k=1,2,\dots,m$, $I=(ij)\in S$:
$v_{i}^{I}(x)=\sqrt{|v^{I}(x)|},\,v_{j}^{I}(x)=\sqrt{|v^{I}(x)|}{\mathrm{sign}}(v^{I}(x)),$ (2.12)
$v_{k}^{I}(x)=0\;\text{for all}\;k\notin\{i,j\},$
$v(x)=-F^{-1}(x)(\nabla V(x))^{\top},$
where the $n\times n$ matrix $F(x)=\left(f_{1}(x),\dots,f_{m}(x),f_{I}(x)_{I\in S}\right)$ is non-singular for all $x\in D$, and $v(x)=(v_{1}^{0}(x),\dots v_{m}^{0}(x),v^{I}(x)_{I\in S})^{\top}$. Assume that there exists a $\gamma>0$ and a closed bounded domain $D^{\prime}\subseteq D$ with $0\in{\mathrm{int}}\,D^{\prime}$ such that the following conditions hold:
  • c1)
    $\sup\limits_{x\in D^{\prime}\setminus\{0\}}\dfrac{\nabla V(x)\Phi(x,\gamma)}{\|\nabla V(x)\|^{2}}<1$, where
    $\Phi(x,\gamma)=\sum\limits_{I=(ij)\in S}\Big{(}(\gamma^{2}-1)[f_{i}^{I},f_{j}^{I}](x)v^{I}(x)+\gamma^{2}f_{j}(x)$
    ×$\times\nabla|v^{I}(x)|f_{i}(x){\mathrm{sign}}(v^{I}(x))-f_{i}(x)\nabla|v^{I}(x)|f_{j}(x)\Big{)};$
  • c2)
    $\sup\limits_{x\in D^{\prime}}\dfrac{|r(x,\varepsilon,\gamma)|}{\|\nabla V(x)\|^{2}}<\infty$ for all $\varepsilon\in(0,\bar{\varepsilon}]$ for some $\bar{\varepsilon}>0$, where $r(x,\varepsilon,\gamma)$ is the remainder in representation (2.8).
Then there exists a $\delta>0$ such that each solution $x(t)$ of system (2.4) with the initial condition $x(0)\in B_{\delta}(0)\subset D^{\prime}$ tends to zero as $t\to+\infty$, provided that the time-varying functions $\phi_{k}^{I,\varepsilon}$ are given by (2.2) under assumption (2.3), and the parameter $\varepsilon>0$ is small enough.
Proof.
The proof is similar to the proof of Theorem 2.1. Under the assumptions of Theorem 2.2, the function $W$ defined by (2.9) can be represented as
$W(x)$ $=\nabla V(x)\Big{(}g_{0}(x)+\gamma^{2}\sum_{I=(ij)\in S}\big{(}[f_{i}^{I},f_{j}^{I}](x)v^{I}(x)$
$+f_{j}(x)$ $\nabla|v^{I}(x)|f_{i}(x){\mathrm{sign}}(v^{I}(x))-f_{i}(x)\nabla|v^{I}(x)|f_{j}(x)\Big{)}$
$=\nabla V(x)F(x)v(x)+\nabla V(x)r_{f}(x)$
$=-\|\nabla V(x)\|^{2}+\nabla V(x)\Phi(x,\gamma).$
By c1), there exist $\gamma>0$ and $\sigma\in[0,1]$, such that
$W(x)\leq-(1-\sigma)\|\nabla V(x)\|^{2}\text{ in }D^{\prime}.$
Additionally, condition c2) implies
$|r(x,\varepsilon,\gamma)|\leq\tilde{c}_{r}\|\nabla V(x)\|^{2}$
in $D^{\prime}$. Thus, representation (2.8) implies
$V(x(\varepsilon))\leq V(x^{0})-\varepsilon\|\nabla V(x^{0})\|^{2}\big{(}1-\sigma-\sqrt{\varepsilon}\tilde{c}_{r}\big{)}.$
If $\varepsilon\in\big{(}0,{\mathrm{min}}\{((1-\sigma)/\tilde{c}_{r})^{2},\bar{\varepsilon}\}\big{)}$, then $V(x(\varepsilon))<V(x^{0})$ for $x(\varepsilon)\neq x^{0}$, and the rest of the proof goes along the same line as the proof of Theorem 2.1.
Under more general assumptions, the condition ii) of Theorem 2.1 can be established by the following technical lemma.
Lemma 2.3.
Let a function $V:D\to\mathbb{R}$ be differentiable such that $\nabla V(0)=0$. Assume that:
  • a)
    $\alpha(x):=L_{g_{0}}V(x)\leq 0$ for all $x\in D$;
  • b)
    $\beta(x):=\sum\limits_{I=(ij)\in S}L_{[g_{i}^{I},g_{j}^{I}]}V(x)<0$ for all $x\in M_{0}\setminus\{0\}$, where $M_{0}=\{x\in D\,|\,\alpha(x)=0\};$
  • c)
    $\sup\limits_{x\in D\setminus M_{0}}\left\{-\dfrac{\beta(x)}{\alpha(x)}\right\}<\infty.$
Then there exists a $\gamma>0$ such that the function $W(x)$ defined by (2.9) is negative definite.
Proof.
Indeed, under assumption b), the function
$W(x)=\alpha(x)+\gamma^{2}\beta(x)$
is strictly negative in $M_{0}\setminus\{0\}$ for each $\gamma\neq 0$, and $W(0)=0$ because $\nabla V(0)=0$. It remains to show that $W(x)<0$ for all $x\in D\setminus M_{0}$. Let $c_{\alpha\beta}=\sup\limits_{x\in D\setminus M_{0}}\left\{-\frac{\beta(x)}{\alpha(x)}\right\}<\infty$. We consider the two possible cases.
Case 1: $c_{\alpha\beta}\leq 0$. Them $W(x)<0$ for all real $\gamma$ and all $x\in D\setminus M_{0}$.
Case 2. $c_{\alpha\beta}>0$. In this case, $W(x)<0$ for all $x\in D\setminus M_{0}$, provided that $\gamma^{2}<1/c_{\alpha\beta}$.
We will illustrate the proposed control design scheme with a multidimensional nonholonomic system in the following section.
3 CASE STUDY
3.1 Multidimensional Brockett Integrator
Consider the following underactuated nonlinear control system:
$\dot{x}=\sum_{j=1}^{4}u_{j}f_{j}(x),\quad x\in{\mathbb{R}}^{10},\;u\in{\mathbb{R}}^{4},$ (3.1)
where $u=(u_{1},u_{2},u_{3},u_{4})^{\top}$, $x=(x_{1},x_{2},\dots,x_{10})^{\top}$,
$f_{1}(x)$ $=(1,0,0,0,-x_{2},-x_{3},-x_{4},0,0,0)^{\top},$
$f_{2}(x)$ $=(0,1,0,0,x_{1},0,0,-x_{3},-x_{4},0)^{\top},$
$f_{3}(x)$ $=(0,0,1,0,0,x_{1},0,x_{2},0,-x_{4})^{\top},$
$f_{4}(x)$ $=(0,0,0,1,0,0,x_{1},0,x_{2},x_{3})^{\top}.$
System (3.1) belongs to the class of nilpotent systems described by R. W. Brockett in [17]. It is easy to show that this system satisfies the controllability rank condition with the Lie brackets $f^{(ij)}(x):=[f_{i},f_{j}](x)$ for $(ij)\in S$,
$S=\{(12),\,(13),\,(14),\,(23),\,(24),\,(34)\}.$
Indeed,
$f^{(12)}(x)=2e_{5},\,f^{(13)}(x)=2e_{6},$
$f^{(14)}(x)=2e_{7},\,f^{(23)}(x)=2e_{8},$
$f^{(24)}(x)=2e_{9},\,f^{(34)}(x)=2e_{10},$
where $e_{j}\in\mathbb{R}^{10}$ denotes the unit column vector with $1$ at the $j$-th entry. Thus,
${\mathrm{span}}(f_{1}(x),\dots,f_{4}(x),f^{I}(x)_{I\in S})={\mathbb{R}}^{10}\;\text{for each}\;x\in\mathbb{R}^{10}.$
System (3.1) is not stabilizable by a continuous time-invariant feedback law of the form $u=h(x)$ because it does not satisfy Brockett’s necessary condition [1]. It should be mentioned that Brockett’s condition remains necessary for a class of partial stabilization problems [18], which poses challenges for stabilizing control design with respect to certain variables. Note that system (3.1) is nilpotent, so that higher-order terms with iterated Lie brackets in the Chen–Fliess expansion of (3.1) vanish.
We take the Lyapunov function candidate as
$V(x)=\frac{1}{2}\sum_{j=1}^{4}x_{j}^{2}+\frac{1}{2m}\sum_{j=5}^{10}x_{j}^{2m},\,m\geq 1,$
and define the following time-varying feedback control in the form (2.1):
$u_{1}{=}v_{1}^{0}(x)+\gamma\sum_{k=2}^{4}v_{1}^{(1k)}(x)\phi_{1}^{(1k),\varepsilon}(t),$ (3.2)
$u_{2}{=}v_{2}^{0}(x)+\gamma\sum_{k=3}^{4}v_{2}^{(2k)}(x)\phi_{2}^{(2k),\varepsilon}(t)+\gamma v_{2}^{(12)}(x)\phi_{2}^{(12),\varepsilon}(t),$
$u_{3}{=}v_{3}^{0}(x)+\gamma\sum_{k=1}^{2}v_{3}^{(k3)}(x)\phi_{3}^{(k3),\varepsilon}(t)+\gamma v_{3}^{(34)}(x)\phi_{3}^{(34),\varepsilon}(t),$
$u_{4}{=}v_{4}^{0}(x)+\gamma\sum_{k=1}^{3}v_{4}^{(k4)}(x)\phi_{4}^{(k4),\varepsilon}(t),$
where
$\phi_{1}^{(12),\varepsilon}(t)=2\sqrt{\frac{\pi}{\varepsilon}}\cos\left(\frac{2\pi t}{\varepsilon}\right),\;\phi_{2}^{(12),\varepsilon}(t)=2\sqrt{\frac{\pi}{\varepsilon}}\sin\left(\frac{2\pi t}{\varepsilon}\right),$
$\phi_{1}^{(13),\varepsilon}(t)=2\sqrt{\frac{2\pi}{\varepsilon}}\cos\left(\frac{4\pi t}{\varepsilon}\right),\;\phi_{3}^{(13),\varepsilon}(t)=2\sqrt{\frac{2\pi}{\varepsilon}}\sin\left(\frac{4\pi t}{\varepsilon}\right),$
$\phi_{1}^{(14),\varepsilon}(t)=2\sqrt{\frac{3\pi}{\varepsilon}}\cos\left(\frac{6\pi t}{\varepsilon}\right),\;\phi_{4}^{(14),\varepsilon}(t)=2\sqrt{\frac{3\pi}{\varepsilon}}\sin\left(\frac{6\pi t}{\varepsilon}\right),$
$\phi_{2}^{(23),\varepsilon}(t)=4\sqrt{\frac{\pi}{\varepsilon}}\cos\left(\frac{8\pi t}{\varepsilon}\right),\;\phi_{3}^{(23),\varepsilon}(t)=4\sqrt{\frac{\pi}{\varepsilon}}\sin\left(\frac{8\pi t}{\varepsilon}\right),$
$\phi_{2}^{(24),\varepsilon}(t)=2\sqrt{\frac{5\pi}{\varepsilon}}\cos\left(\frac{10\pi t}{\varepsilon}\right),\;\phi_{4}^{(24),\varepsilon}(t)=2\sqrt{\frac{5\pi}{\varepsilon}}\sin\left(\frac{10\pi t}{\varepsilon}\right),$
$\phi_{3}^{(34),\varepsilon}(t)=2\sqrt{\frac{6\pi}{\varepsilon}}\cos\left(\frac{12\pi t}{\varepsilon}\right),\;\phi_{4}^{(34),\varepsilon}(t)=2\sqrt{\frac{6\pi}{\varepsilon}}\sin\left(\frac{12\pi t}{\varepsilon}\right),$
$v_{k}^{0}(x)=-x_{k},\quad k\in\{1,2,3,4\}.$
The resulting closed-loop system (3.1), (3.2) takes the form
$\dot{x}=g_{0}(x)+\gamma\sum_{I\in S}\sum_{k=1}^{4}g_{k}^{I}(x)\phi_{k}^{I,\varepsilon}(t),$ (3.3)
with $g_{0}(x)=-(x_{1},x_{2},x_{3},x_{4},0,0,0,0,0,0)^{\top},$ and $g_{k}^{I}(x)=v_{k}^{I}(x)f_{k}(x),\;k\in\{1,2,3,4\}$.
Figure 1: Solution components of the closed-loop system (3.1), (3.2), (3.4).
The solution $x(t)$ of system (3.3) with the initial data $x(0)=x^{0}\in{\mathbb{R}}^{3}$ is represented by the Chen–Fliess series (2.6):
$x(\varepsilon)=x^{0}+\varepsilon\left(g_{0}(x^{0})+{\gamma}^{2}\sum_{I=(ij)\in S}[g_{i}^{I},g_{j}^{I}](x^{0})\right)+O(\varepsilon^{3/2}).$
It is easy to see that $V(x)$ is a weak CLF, but not a strict CLF for system (3.1). According to the general scheme presented in Section 2, we design the functions $v_{j}^{I}(x)$ such that
$W(x)=L_{g_{0}}V(x)+{\gamma}^{2}L_{[g_{1},g_{2}]}V(x)$
is negative definite. Note that
$\alpha(x)=L_{g_{0}}V(x)=-(x_{1}^{2}+x_{2}^{2}+x_{3}^{2}+x_{4}^{2})\leq 0,$
so that condition a) of Lemma 2.3 is satisfied. Let us define
$v_{i}^{I}(x)=\sqrt{|v^{I}(x)|},\,v_{j}^{I}(x)=\sqrt{|v^{I}(x)|}{\mathrm{sign}}(v^{I}(x)),$ (3.4)
$v^{(12)}(x)=-\frac{{\mathrm{sign}}(x_{5})}{2}|x_{5}|^{2m-1},$
$v^{(13)}(x)=-\frac{{\mathrm{sign}}(x_{6})}{2}|x_{6}|^{2m-1},$
$v^{(14)}(x)=-\frac{{\mathrm{sign}}(x_{7})}{2}|x_{7}|^{2m-1},$
$v^{(23)}(x)=-\frac{{\mathrm{sign}}(x_{8})}{2}|x_{8}|^{2m-1},$
$v^{(24)}(x)=-\frac{{\mathrm{sign}}(x_{9})}{2}|x_{9}|^{2m-1},$
$v^{(34)}(x)=-\frac{{\mathrm{sign}}(x_{10})}{2}|x_{10}|^{2m-1}.$
It is easy to see that the above functions $v_{i}^{I}(x)$ and $v_{j}^{I}(x)$ are sufficiently many times continuously differentiable under a suitable choice of $m$. Moreover, according to the notation in Lemma 2.3, $\beta(x)=-\|\tilde{x}\|^{2m}+\dfrac{2m-1}{4}\Phi(x),$ where $\tilde{x}=(x_{5},x_{6},x_{7},x_{8},x_{9},x_{10})^{\top}$,
$\Phi(x)$ $=x_{1}^{2}(|x_{5}|^{2m-2}+|x_{6}|^{2m-2}+|x_{7}|^{2m-2})$
$+x_{2}^{2}(|x_{5}|^{2m-2}+|x_{8}|^{2m-2}+|x_{9}|^{2m-2})$
$+x_{3}^{2}(|x_{6}|^{2m-2}+|x_{8}|^{2m-2}+|x_{10}|^{2m-2})$
$+x_{4}^{2}(|x_{7}|^{2m-2}+|x_{9}|^{2m-2}+|x_{10}|^{2m-2})$
$-|x_{5}|^{2m-2}(x_{1}x_{3}x_{6}+x_{1}x_{4}x_{7}+x_{2}x_{3}x_{8}+x_{2}x_{4}x_{9})$
$-|x_{6}|^{2m-2}(x_{1}x_{2}x_{5}+x_{1}x_{4}x_{7}-x_{2}x_{3}x_{8}+x_{3}x_{4}x_{10})$
$-|x_{7}|^{2m-2}(x_{1}x_{2}x_{5}+x_{1}x_{3}x_{6}-x_{2}x_{4}x_{9}-x_{2}x_{4}x_{10})$
$+|x_{8}|^{2m-2}(x_{1}x_{2}x_{5}+x_{1}x_{3}x_{6}-x_{2}x_{4}x_{9}-x_{3}x_{4}x_{10})$
$+|x_{9}|^{2m-2}(x_{1}x_{2}x_{5}+x_{1}x_{4}x_{7}-x_{2}x_{3}x_{8}+x_{3}x_{4}x_{10})$
$+|x_{10}|^{2m-2}(x_{1}x_{3}x_{6}+x_{1}x_{4}x_{7}+x_{2}x_{3}x_{8}+x_{2}x_{4}x_{9}).$
Thus, for all $x\in\mathbb{R}^{10}$,
$\|\Phi(x)\|\leq(x_{1}^{2}+x_{2}^{2}+x_{3}^{2}+x_{4}^{2})\|\tilde{x}\|^{2m-2}(1+\|\tilde{x}\|),$
and
$W(x)=-$ $(x_{1}^{2}+x_{2}^{2}+x_{3}^{2}+x_{4}^{2})-\gamma^{2}\|\tilde{x}\|^{2m}$ (3.5)
$+\frac{(2m-1)\gamma^{2}}{4}\Phi(x)$
$\leq$-≤ - $(x_{1}^{2}+x_{2}^{2}+x_{3}^{2}+x_{4}^{2})$
×$\times\Big{(}1-\frac{(2m-1)\gamma^{2}}{4}$ $\|\tilde{x}\|^{2m-2}(1+\|\tilde{x}\|)\Big{)}-\gamma^{2}\|\tilde{x}\|^{2m}.$
The formal derivation of $W(x)$ is valid for $x_{j}\neq 0$, $j\in\{5,6,\dots,10\}$, and $m\in[1,\frac{3}{2}]$, and for any $x_{j}$ if $m>\frac{3}{2}$. Note that the resulting expression (3.5) is well-defined for all $x\in{\mathbb{R}}^{10}$ and $m\geq\frac{3}{2}$. It can be shown that condition b) of Lemma 2.3 is satisfied in any bounded domain $D$ if $m<\frac{5}{4}$, and condition ii) holds for $D=\mathbb{R}^{10}$ if $m\geq\frac{5}{4}$. Thus, the following corollary of Theorem 2.1 summarizes this case study: there exist $\varepsilon>0$, $\gamma>0$, and $\delta>0$ such that each solution $x(t)$ of the closed-loop system (3.1), (3.2), (3.4) with $x(0)\in B_{\delta}(0)$ tends to zero as $t\to+\infty$.
In the considered example, it is easy to see how to define the gain parameters to ensure that $W(x)$ is negative definite. In particular, $W(x)$ is negative definite in the domain $D_{H}=\{x\in\mathbb{R}\,|\;\|\tilde{x}\|<H\}$, $H>0$, if the following conditions are satisfied:
$m>1,\;\gamma<\left(0,\frac{2}{\sqrt{(2m-1)H^{2m-1}(1+H)}}\right),$ (3.6)
Moreover, $W(x)$ is negative definite in ${\mathbb{R}}^{10}$, provided that
$m=1,\;\gamma\in\left(0,\sqrt{2}\right).$ (3.7)
To illustrate that the proposed feedback control ensures asymptotic stability of the multidimensional Brockett integrator, we perform numerical integration of the closed-loop system (3.1), (3.2), (3.4) under conditions (3.6) and (3.7). Fig. 1 illustrates the behavior of the closed-loop system with $m=\frac{1}{2}$ (left) and $m=1$ (right). As one can see, in both cases, the solutions tend to zero for large $t$, while the norm $\|x(t)\|$ exhibits exponential convergence to 00 if $m=1$, and a slower (polynomial) convergence if $m=\frac{3}{2}$. In both cases, we set $\varepsilon=0.1$ and $\gamma=0.5$. The initial conditions are chosen as
$x(0)=\left(1,-1,1.5,-0.5,2,-2,2.5,-2.5,3,-3\right)^{\top},$
in the first case, and
$x(0)=\left(1,-1,1,-1,1,-1,1,-1,1,-1\right)^{\top}$
in the second one.
4 CONCLUSIONS AND FUTURE WORK
Our main theoretical contribution is formulated in Theorem 2.1 and illustrated by the multidimensional Brockett integrator. One of the key requirements in Theorem 2.1 and Theorem 2.2 is the $C^{2}$ assumption on the control functions and the relevant vector fields $g_{0}$ and $g_{k}^{I}$. As it can be seen in Section 3, this regularity property holds if the exponent $m$, appearing in the Lyapunov function candidate, is suitably chosen.
We expect to relax this regularity assumption while still ensuring practical stabilizability, meaning attraction to a neighborhood of the origin, where the neighborhood’s radius decreases as $\varepsilon$ decreases. This result can be established similarly to the proof of practical stability in [19, Theorem 3, Assertion I]. Additionally, we intend to explore the problem of practical stabilization when condition iii) of Theorem 2.1 is violated. In future work, we also plan to develop a systematic approach for constructing the functions $v^{I}_{k}(x)$ in (2.1), based on the conditions on $\beta(x)$ of Lemma 2.3.